# Bayesian Additive Regression Trees for functional ANOVA model

### v2

@@ D.3 Rademacher complexity bound
E.4 Proof of (E.4)
For given $\mathbf{x}^{(n)}\in A_{n}$, where $A_{n}$ is defined in Section C.3 of Supplementary Material, we have
${\pi_{\xi}\{\Theta_{\xi}\backslash\Theta_{\xi}^{n}\}\over\pi_{\xi}\{\mathbb{B}_{n}\}}$ $\leq\exp\left(-2n\epsilon_{n}^{2}\right)$
since Condition (C.17) and Condition (C.18) hold. Thus, Lemma 1 of nonp1 implies that for any $\mathbf{x}^{(n)}\in A_{n}$ and an arbitrary $\delta>0$,
$\lim_{n\to\infty}\mathbb{P}_{Y^{(n)}}\bigg{\{}\pi_{\xi}\{\Theta_{\xi}\backslash\Theta_{\xi}^{n}|\mathbf{X}^{(n)},Y^{(n)}\}>\delta\bigg{|}\mathbf{X}^{(n)}=\mathbf{x}^{(n)}\bigg{\}}=0.$
@@
$\lim_{n\to\infty}\mathbb{P}_{0}^{n}\bigg{\{}\pi_{\xi}\{\Theta_{\xi}\backslash\Theta_{\xi}^{n}|\mathbf{X}^{(n)},Y^{(n)}\}>\delta\bigg{\}}=0,$
for any $\delta>0$, which completes the proof of (E.4). ∎
Appendix F Proof of Theorem C.1
The proof consists of three steps. In the first step, we approximate $f_{0,S}$ by a specially designed multinary-product tree called the equal probability-product tree, which is populationally identifiable. The second step is to approximate the equal probability-product tree by an identifiable multinary-product tree using Theorem D.2. The third step is to transfer the identifiable multinary-product tree obtained in the second step into a sum of identifiable binary-product trees whose heights are bounded.
F.1 Approximation of $f_{0,S}$ by the equal probability-product tree
@@
F.2 Approximation of the EP-product tree by an identifiable multinary-product tree
Since the EP-product tree $f_{S,\mathcal{P}_{r}^{EP},\gamma}$ in Section F.1 is a populationally identifiable multinary-product tree, we can apply Theorem D.2 to obtain an identifiable multinary-product tree $f_{S,\mathcal{P}_{r}^{EP},\hat{\gamma}}$ such that
@@ F.2 Approximation of the EP-product tree by an identifiable multinary-product tree
$\leq\|f_{0,S}\|_{\mathcal{H}^{\alpha}}\bigg{(}{m\over r}\bigg{)}^{\alpha}+(24F+1)2^{|S|}\sqrt{r^{|S|}\log n\over n}.$
with probability at least $1-2^{|S|+1}|S|r^{|S|^{2}}\exp\left(-{r^{|S|}\log n\over 4F^{2}}\right)$ with resepect to $\mathbb{P}^{n}_{\mathbf{X}}$, where $m=\sqrt{|S|}/p_{L}.$
F.3 Decomposition of the modified EP-product tree by a sum of identifiable binary-product trees with bounded heights
@@ F.3.1 Notations
Figure F.4: Example of decomposing the partition of $f$ into the partitions of $f_{1}^{(1)}$ and $f_{2}^{(1)}$.
F.3.2 Proof of the decomposition
The strategy of the decomposition is to decompose a given identifiable multinary-product tree into the sum of two sibling identifiable multinary-product trees whose orders are smaller than their parent identifiable multinary-product tree. Figure F.4 presents an example of decomposing the partition of an identifiable multinary-product tree $f$ for $S=\{1,2\}$ and $\phi_{1}=\phi_{2}=3$ into two partitions of child identifiable multinary-product trees $f_{1}^{(1)}$ and $f_{2}^{(1)}$, where $\text{order}(f_{1})_{1}=2$, $\text{order}(f_{1})_{2}=3$, $\text{order}(f_{2})_{1}=2$, and $\text{order}(f_{2})_{2}=3$. We repeat this decomposition until the orders of all identifiable multinary-product trees become 2. Figure F.5 presents an example of decomposing the partition of a identifiable multinary-product tree $f$ into four partitions of identifiable binary-product trees $f_{1,1}^{(1,2)},f_{1,2}^{(1,2)},f_{2,1}^{(1,2)}$, and $f_{2,2}^{(1,2)}$, where $\text{order}(f_{i,j}^{(1,2)})_{k}=2$ for $i=1,2$, $k=1,2$ and $j=1,2$. The following lemma is the key tool whose proof is given in Section F.3.3 and F.3.4.
Lemma F.2.
@@ F.3.2 Proof of the decomposition
Figure F.5: Example of decomposing partition of $f$ into the partitions of identifiable binary-product trees $f_{1,1}^{(1,2)}$, $f_{1,2}^{(1,2)}$, $f_{2,1}^{(1,2)}$, and $f_{2,2}^{(1,2)}$.
We will prove the decomposition by use of Lemma F.2. Let $f$ be a given EP-product tree of the component $S$ with ${\rm order}(f)_{j}=r$ for all $j\in S,$ which we are going to decompose. Without loss of generality, we let $S=\{1,2,\ldots,|S|\}.$ At first, we apply Lemma F.2 to decompose $f(\cdot)=\sum_{k_{1}=1}^{r-1}f_{k_{1}}^{(1)}(\cdot),$ where $f_{k_{1}}^{(1)}$s are identifiable multinary-product trees such that $\text{order}(f_{k_{1}}^{(1)})_{1}=2$ and $\text{order}(f_{k_{1}}^{(1)})_{\ell}=r,\ell\geq 2$ for all $k_{1}=1,\ldots,r-1$ and $\sup_{k_{1}}\|f_{k_{1}}^{(1)}\|_{\infty}\leq 2\|f\|_{\infty}.$
@@
$\textbf{u}=\sum_{j=1}^{r-1}(a_{j}\mathbb{1}_{j}+b_{j}\mathbb{1}_{j}^{c}),$ (F.1)
$\mathbb{w}^{\top}(a_{j}\mathbb{1}_{j}+b_{j}\mathbb{1}_{j}^{c})=0$ (F.2)
@@
Figure F.6: Decomposition of $\mathbf{u}^{(0)}$ into $\mathbf{u}^{(1)}$ and $(a_{1}\mathbb{1}_{1}+b_{1}\mathbb{1}_{1}^{c})$ in the case of $r=8$.
@@
where $a_{1}=(1-w_{1})(u_{1}^{(0)}-u_{2}^{(0)})$ and $b_{1}=w_{1}(u_{2}^{(0)}-u_{1}^{(0)}).$ It is easy to see that $\mathbb{w}^{\top}(a_{1}\mathbb{1}_{1}+b_{1}\mathbb{1}_{1}^{c})=0$ and thus $\mathbb{w}^{\top}\textbf{u}^{(1)}=0$. Moreover, we have $u_{j}^{(1)}=u_{j}^{(0)}-b_{1}$ for $j=3,\ldots,r$ and $u_{1}^{(1)}=u_{1}^{(0)}-a_{1}.$ In addition, it can be shown that $u_{1}^{(1)}=u_{2}^{(1)}.$ Figure F.6 presents an example of this decomposition in the case of $r=8$. For $k\in\{2,\ldots,r-2\},$ we define $\textbf{u}^{(k)}$ recursively as
$\textbf{u}^{(k)}:=\textbf{u}^{(k-1)}-(a_{k}\mathbb{1}_{k}+b_{k}\mathbb{1}_{k}^{c}),$ (F.3)
@@
Here, (F.5), (F.6), and (F.7) are due to $u_{1}^{(r-3)}=\cdots=u_{r-2}^{(r-3)}$ and (F.4). Thus, (F.3) holds with $\textbf{u}^{(r-1)}={\bf 0}.$ Hence, we complete the proof of the existence of the real numbers $(a_{j},b_{j}),j=1,...,r-1$, satisfies (F.1).
@@
F.3.4 Proof of Lemma F.2 for general $S$
Fix $\boldsymbol{\ell}\in{\rm index}(f)$ and $h\in S$. Let $\textbf{u}_{\boldsymbol{\ell},h}=(\gamma_{\boldsymbol{\ell}_{+(h,j)}},j\in[r]).$ Applying Claim in Section F.3.3 of Supplementary Material to $\textbf{u}_{\boldsymbol{\ell},h}$ with $\mathbb{w}=(\mu_{n,h}\{I_{h,j}\},j\in[r])$, we have
@@
for $k\in S$. First, $\int_{\mathcal{X}_{k}}f_{j}^{(h)}(\textbf{x}_{S})\mu_{n,k}(dx_{k})=0$ for $k=h$ by (F.2). For $k\neq h,$ the proof of Claim reveals that there exists $r$-dimensional vectors $\mathbf{v}_{1}^{(j)}$ and $\mathbf{v}_{2}^{(j)}$ such that $\gamma^{(j)}_{\boldsymbol{\ell}_{+(h,m)}}=\mathbf{v}_{m}^{(j)\top}\gamma_{\boldsymbol{\ell}_{+(h,\cdot)}},$ for $m=1,2,$ where $\gamma_{\boldsymbol{\ell}_{+(h,\cdot)}}=(\gamma_{\boldsymbol{\ell}_{+(h,j)}},j\in[r]).$ Thus, for $m=1,2$, we have
$\mathbb{E}_{\mu_{n},k}(\gamma_{\boldsymbol{\ell}_{+(h,m)}}^{(j)})=\mathbb{v}_{m}^{(j)\top}\mathbb{E}_{\mu_{n},k}(\gamma_{\boldsymbol{\ell}_{+(h,\cdot)}})=0$ (F.8)
@@
and $\mathbb{E}_{\mu_{n},k}(\cdot)$ is defined in (D.1).
@@ Appendix G Discussion about the posterior concentration rate for the $X$-fixed Design
A key issue of ANOVA-BART for the $X$-fixed design is how to define $f_{0,S}$. For a given $f_{0}$, Theorem 2.1 yields a unique ANOVA decomposition with respect to $\mu_{n}$, whose interactions are identifiable. Recall that the proof of the posterior convergence rate for the $X$-random design consists of the two main parts: (1) for $\mathbf{x}^{(n)}\in A_{n},$ where $A_{n}$ is defined in (C.8), the posterior convergence rate is $\epsilon_{n}$ and (2) the probability of $A_{n}$ converges to 1. Thus, for the $X$-fixed design, the mission is to figure out sufficient conditions for $\mathbf{x}^{(n)}\in A_{n}$ with $f_{0,S}$ defined in Theorem 2.1 with respect to $\mu_{n}$.
In order that $\mathbf{x}^{(n)}\in A_{n},$ there exists a sum of identifiable binary-product trees that approximates $f_{0,S}$ closely for each $S\subseteq[p].$ In Section F.1, we show that the EP-tree, which satisfies the populational identifiability condition, approximates $f_{0,S}$ closely for the $X$-random design. We can modify this proof to show that the empirical EP-tree, which satisfies the identifability condition, approximates $f_{0,S}$ closely for the $X$-fixed design when there exists a positive constant $C_{2}$ such that, for all $r\in[n],$
@@
where ${\rm Diam}(A):=\max_{\mathbf{v},\mathbf{w}\in A}\|\mathbf{v}-\mathbf{w}\|_{2}$ and $\mathcal{P}_{r}^{nEP}$ denotes the equal-probability partition induced by the quantiles of the empirical distribution. Then, by letting the modified EP-tree defined in Section F.2 to be equal to the empirical EP-tree, all the results in Section F.2 to Section F.3 are satisfied. Thus, we conclude that the posterior convergence rate is $\epsilon_{n}$ as long as $\max_{A\in\mathcal{P}_{r}^{nEP}}\text{Diam}(A)\leq C_{2}/r$ for all $r\in[n].$

### v3

@@ C.3 Rademacher complexity bound
D.4 Proof of (D.4)
For given $\mathbf{x}^{(n)}\in A_{n}$, where $A_{n}$ is defined in Section B.3 of Supplementary Material, we have
${\pi_{\xi}\{\mathcal{F}_{\xi}\backslash\mathcal{F}_{\xi}^{n}\}\over\pi_{\xi}\{\mathbb{B}_{n}\}}$ $\leq\exp\left(-2n\epsilon_{n}^{2}\right)$
since Condition (B.19) and Condition (B.20) hold. Thus, Lemma 1 of Ghosal and van der Vaart (2007) implies that for any $\mathbf{x}^{(n)}\in A_{n}$ and an arbitrary $\delta>0$,
$\lim_{n\to\infty}\mathbb{P}_{Y^{(n)}}\bigg\{\pi_{\xi}\{\mathcal{F}_{\xi}\backslash\mathcal{F}_{\xi}^{n}|\mathbf{X}^{(n)},Y^{(n)}\}>\delta\bigg|\mathbf{X}^{(n)}=\mathbf{x}^{(n)}\bigg\}=0.$
@@
$\lim_{n\to\infty}\mathbb{P}_{0}^{n}\bigg\{\pi_{\xi}\{\mathcal{F}_{\xi}\backslash\mathcal{F}_{\xi}^{n}|\mathbf{X}^{(n)},Y^{(n)}\}>\delta\bigg\}=0,$
for any $\delta>0$, which completes the proof of (D.4). ∎
Appendix E Proof of Theorem B.1
The proof consists of three steps. In the first step, we approximate $f_{0,S}$ by a specially designed multinary-product tree called the equal probability-product tree, which is populationally identifiable. The second step is to approximate the equal probability-product tree by an identifiable multinary-product tree using Theorem C.2. The third step is to transfer the identifiable multinary-product tree obtained in the second step into a sum of identifiable binary-product trees whose heights are bounded.
E.1 Approximation of $f_{0,S}$ by the equal probability-product tree
@@
E.2 Approximation of the EP-product tree by an identifiable multinary-product tree
Since the EP-product tree $f_{S,\mathcal{P}_{r}^{EP},\gamma}$ in Section E.1 is a populationally identifiable multinary-product tree, we can apply Theorem C.2 to obtain an identifiable multinary-product tree $f_{S,\mathcal{P}_{r}^{EP},\hat{\gamma}}$ such that
@@ E.2 Approximation of the EP-product tree by an identifiable multinary-product tree
$\leq\|f_{0,S}\|_{\mathcal{H}^{\alpha}}\bigg({C_{S}\over r}\bigg)^{\alpha}+(24F+1)2^{|S|}\sqrt{r^{|S|}\log n\over n}.$
with probability at least $1-2^{|S|+1}|S|r^{|S|^{2}}\exp\left(-{r^{|S|}\log n\over 4F^{2}}\right)$ with resepect to $\mathbb{P}^{n}_{\mathbf{X}}$, where $C_{S}=\sqrt{|S|}/p_{L}$.
E.3 Decomposition of the modified EP-product tree by a sum of identifiable binary-product trees with bounded heights
@@ E.3.1 Notations
Figure E.5: Example of decomposing the partition of $f$ into the partitions of $f_{1}^{(1)}$ and $f_{2}^{(1)}$.
E.3.2 Proof of the decomposition
The strategy of the decomposition is to decompose a given identifiable multinary-product tree into the sum of two sibling identifiable multinary-product trees whose orders are smaller than their parent identifiable multinary-product tree. Figure E.5 presents an example of decomposing the partition of an identifiable multinary-product tree $f$ for $S=\{1,2\}$ and $\phi_{1}=\phi_{2}=3$ into two partitions of child identifiable multinary-product trees $f_{1}^{(1)}$ and $f_{2}^{(1)}$, where $\text{order}(f_{1})_{1}=2$, $\text{order}(f_{1})_{2}=3$, $\text{order}(f_{2})_{1}=2$, and $\text{order}(f_{2})_{2}=3$. We repeat this decomposition until the orders of all identifiable multinary-product trees become 2. Figure E.6 presents an example of decomposing the partition of a identifiable multinary-product tree $f$ into four partitions of identifiable binary-product trees $f_{1,1}^{(1,2)},f_{1,2}^{(1,2)},f_{2,1}^{(1,2)}$, and $f_{2,2}^{(1,2)}$, where $\text{order}(f_{i,j}^{(1,2)})_{k}=2$ for $i=1,2$, $k=1,2$ and $j=1,2$. The following lemma is the key tool whose proof is given in Section E.3.3 and E.3.4.
Lemma E.2.
@@ E.3.2 Proof of the decomposition
Figure E.6: Example of decomposing partition of $f$ into the partitions of identifiable binary-product trees $f_{1,1}^{(1,2)}$, $f_{1,2}^{(1,2)}$, $f_{2,1}^{(1,2)}$, and $f_{2,2}^{(1,2)}$.
We will prove the decomposition by use of Lemma E.2. Let $f$ be a given EP-product tree of the component $S$ with ${\rm order}(f)_{j}=r$ for all $j\in S,$ which we are going to decompose. Without loss of generality, we let $S=\{1,2,\ldots,|S|\}.$ At first, we apply Lemma E.2 to decompose $f(\cdot)=\sum_{k_{1}=1}^{r-1}f_{k_{1}}^{(1)}(\cdot),$ where $f_{k_{1}}^{(1)}$s are identifiable multinary-product trees such that $\text{order}(f_{k_{1}}^{(1)})_{1}=2$ and $\text{order}(f_{k_{1}}^{(1)})_{\ell}=r,\ell\geq 2$ for all $k_{1}=1,\ldots,r-1$ and $\sup_{k_{1}}\|f_{k_{1}}^{(1)}\|_{\infty}\leq 2\|f\|_{\infty}.$
@@
$\textbf{u}=\sum_{j=1}^{r-1}(a_{j}\mathbb{1}_{j}+b_{j}\mathbb{1}_{j}^{c}),$ (E.1)
$\mathbb{w}^{\top}(a_{j}\mathbb{1}_{j}+b_{j}\mathbb{1}_{j}^{c})=0$ (E.2)
@@
Figure E.7: Decomposition of $\mathbf{u}^{(0)}$ into $\mathbf{u}^{(1)}$ and $(a_{1}\mathbb{1}_{1}+b_{1}\mathbb{1}_{1}^{c})$ in the case of $r=8$.
@@
where $a_{1}=(1-w_{1})(u_{1}^{(0)}-u_{2}^{(0)})$ and $b_{1}=w_{1}(u_{2}^{(0)}-u_{1}^{(0)}).$ It is easy to see that $\mathbb{w}^{\top}(a_{1}\mathbb{1}_{1}+b_{1}\mathbb{1}_{1}^{c})=0$ and thus $\mathbb{w}^{\top}\textbf{u}^{(1)}=0$. Moreover, we have $u_{j}^{(1)}=u_{j}^{(0)}-b_{1}$ for $j=3,\ldots,r$ and $u_{1}^{(1)}=u_{1}^{(0)}-a_{1}.$ In addition, it can be shown that $u_{1}^{(1)}=u_{2}^{(1)}.$ Figure E.7 presents an example of this decomposition in the case of $r=8$. For $k\in\{2,\ldots,r-2\},$ we define $\textbf{u}^{(k)}$ recursively as
$\textbf{u}^{(k)}:=\textbf{u}^{(k-1)}-(a_{k}\mathbb{1}_{k}+b_{k}\mathbb{1}_{k}^{c}),$ (E.3)
@@
Here, (E.5), (E.6), and (E.7) are due to $u_{1}^{(r-3)}=\cdots=u_{r-2}^{(r-3)}$ and (E.4). Thus, (E.3) holds with $\textbf{u}^{(r-1)}={\bf 0}.$ Hence, we complete the proof of the existence of the real numbers $(a_{j},b_{j}),j=1,...,r-1$, satisfies (E.1).
@@
E.3.4 Proof of Lemma E.2 for general $S$
Fix $\boldsymbol{\ell}\in{\rm index}(f)$ and $h\in S$. Let $\textbf{u}_{\boldsymbol{\ell},h}=(\gamma_{\boldsymbol{\ell}_{+(h,j)}},j\in[r]).$ Applying Claim in Section E.3.3 of Supplementary Material to $\textbf{u}_{\boldsymbol{\ell},h}$ with $\mathbb{w}=(\mu_{n,h}\{I_{h,j}\},j\in[r])$, we have
@@
for $k\in S$. First, $\int_{\mathcal{X}_{k}}f_{j}^{(h)}(\textbf{x}_{S})\mu_{n,k}(dx_{k})=0$ for $k=h$ by (E.2). For $k\neq h,$ the proof of Claim reveals that there exists $r$-dimensional vectors $\mathbf{v}_{1}^{(j)}$ and $\mathbf{v}_{2}^{(j)}$ such that $\gamma^{(j)}_{\boldsymbol{\ell}_{+(h,m)}}=\mathbf{v}_{m}^{(j)\top}\gamma_{\boldsymbol{\ell}_{+(h,\cdot)}},$ for $m=1,2,$ where $\gamma_{\boldsymbol{\ell}_{+(h,\cdot)}}=(\gamma_{\boldsymbol{\ell}_{+(h,j)}},j\in[r]).$ Thus, for $m=1,2$, we have
$\mathbb{E}_{\mu_{n},k}(\gamma_{\boldsymbol{\ell}_{+(h,m)}}^{(j)})=\mathbb{v}_{m}^{(j)\top}\mathbb{E}_{\mu_{n},k}(\gamma_{\boldsymbol{\ell}_{+(h,\cdot)}})=0$ (E.8)
@@
and $\mathbb{E}_{\mu_{n},k}(\cdot)$ is defined in (C.1).
@@
Appendix F Discussion about the posterior concentration rate in the Gaussian regression model with unknown nuisance parameter $\sigma^{2}$
In this section, we discuss the posterior concentration rate of ANOVA-BART when the data distribution follows the Gaussian regression model with unknown variance $\sigma^{2}$. That is, we consider the nonparametric regression model defined as
$Y=f(\mathbf{x})+\epsilon,$ (E.1)
where $\epsilon\sim N(0,\sigma^{2})$ and $f$ is a regression function. Among the regularity Conditions (J.1)-(J.4), we replace Condition (J.3) by Condition (K. 1) in place of Condition (J.3) given as
  1. (K. 1)
    There exist positive constants $\sigma_{\text{min}}^{2}$ and $\sigma_{\text{max}}^{2}$ such that $\sigma_{\text{min}}^{2}<\sigma^{2}<\sigma_{\text{max}}^{2}$.
For $\xi>\max\{2^{p}F,\sigma_{\text{min}}^{-2},\sigma_{\text{max}}^{2}\}$, we let $\pi_{\xi}\{\cdot\}\propto\pi\{\}\mathbb{I}(\|f\|_{\infty}\leq\xi,1/\xi\leq\sigma^{2}\leq\xi)$. Using Lemma 1 in Lim and Jeong (2023) and Theorem 4 of Ghosal and van der Vaart (2007), to show the posterior concentration rate of ANOVA-BART in the Gaussian regression model, it suffices to verify the three conditions in (B.18), (B.19) and (B.20), as in the proof of Theorem 2 in Jeong and Ročlková (2023). Since the only part that differs from the existing proof for exponential families is Condition (B.19), we provide a brief explanation of the proof of Condition (B.19) only. Specifically, for Kullback-Leibler ball $\mathbb{B}_{n}$ with parameter $\theta=(f,\sigma^{2})$ and $\theta_{0}=(f_{0},\sigma_{0}^{2})$, direct calculation yields,
$\mathbb{B}_{n}\supseteq\{\theta\in\mathcal{F}_{\xi}\times(1/\xi,\xi):\|f-f_{0}\|_{2,n}\leq C_{\mathbb{B}^{*}}\epsilon_{n},|\sigma^{2}-\sigma_{0}^{2}|\leq C_{\mathbb{B}^{*}}\epsilon_{n}\}$ (E.2)
for some positive constant $C_{\mathbb{B}^{*}}$. Since
$\pi_{\xi}\{\mathbb{B}_{n}\}$ $\geq\pi_{\xi}\{\theta\in\mathcal{F}_{\xi}\times(1/\xi,\xi):\|f-f_{0}\|_{2,n}\leq C_{\mathbb{B}^{*}}\epsilon_{n},|\sigma^{2}-\sigma_{0}^{2}|\leq C_{\mathbb{B}^{*}}\epsilon_{n}\}$ (E.3)
$\geq\pi\{f:\|f-f_{0}\|_{\infty}\leq C_{\mathbb{B}^{*}}\epsilon_{n}\}\pi\{\sigma^{2}:|\sigma^{2}-\sigma^{2}_{0}|\leq C_{\mathbb{B}^{*}}\epsilon_{n}\}$ (E.4)
and
$\pi\bigl\{\sigma^{2}:\lvert\sigma^{2}-\sigma_{0}^{2}\rvert\leq\epsilon_{n}\bigr\}\gtrsim n^{-1},$
Condition (B.19) is satisfied for any $\xi>\max\bigl\{2^{p}F,\ \sigma_{\min}^{-2},\ \sigma_{\max}^{2}\bigr\}$.
@@ Appendix G Discussion about the posterior concentration rate for the $X$-fixed Design
A key issue of ANOVA-BART for the $X$-fixed design is how to define $f_{0,S}$. For a given $f_{0}$, Theorem 2.1 yields a unique ANOVA decomposition with respect to $\mu_{n}$, whose interactions are identifiable. Recall that the proof of the posterior convergence rate for the $X$-random design consists of the two main parts: (1) for $\mathbf{x}^{(n)}\in A_{n},$ where $A_{n}$ is defined in (B.8), the posterior convergence rate is $\epsilon_{n}$ and (2) the probability of $A_{n}$ converges to 1. Thus, for the $X$-fixed design, the mission is to figure out sufficient conditions for $\mathbf{x}^{(n)}\in A_{n}$ with $f_{0,S}$ defined in Theorem 2.1 with respect to $\mu_{n}$.
In order that $\mathbf{x}^{(n)}\in A_{n},$ there exists a sum of identifiable binary-product trees that approximates $f_{0,S}$ closely for each $S\in\mathbb{S}.$ In Section E.1, we show that the EP-tree, which satisfies the populational identifiability condition, approximates $f_{0,S}$ closely for the $X$-random design. We can modify this proof to show that the empirical EP-tree, which satisfies the identifability condition, approximates $f_{0,S}$ closely for the $X$-fixed design when there exists a positive constant $C_{2}$ such that, for all $r\in[n],$
@@
where ${\rm Diam}(A):=\max_{\mathbf{v},\mathbf{w}\in A}\|\mathbf{v}-\mathbf{w}\|_{2}$ and $\mathcal{P}_{r}^{nEP}$ denotes the equal-probability partition induced by the quantiles of the empirical distribution. Then, by letting the modified EP-tree defined in Section E.2 to be equal to the empirical EP-tree, all the results in Section E.2 to Section E.3 are satisfied. Thus, we conclude that the posterior convergence rate is $\epsilon_{n}$ as long as $\max_{A\in\mathcal{P}_{r}^{nEP}}\text{Diam}(A)\leq C_{2}/r$ for all $r\in[n].$